# Pretraining on the Test Set Is No Longer All You Need: A Debate-Driven Approach to QA Benchmarks

### v1

@@ 1 Introduction
We introduce a systematic, reproducible evaluation pipeline transforming standard QA datasets into structured adversarial debates, effectively reducing subjective biases, clearly signaling model reasoning capabilities, and significantly decreasing benchmark curation costs. Empirical experiments with various commercial and open-source models on MMLU-Pro (Wang et al., 2024) validate our method’s robustness and resistance to data contamination. Additionally, weaker judges reliably distinguished stronger debaters, reinforcing findings by Khan et al. (2024), thereby ensuring scalability and future-proof applicability for advanced AI evaluations. We plan to publicly release a standardized benchmark derived from 5,500 structured debates—totaling over 11,000 rounds of argumentation—involving eleven reference models, providing the research community with a dependable evaluation resource. 111Code for this project is available at: github.com/l6cao/Debate-Driven-Evaluation.
@@ 4 Experimental Setup
To assess data contamination effects, we performed a fine-tuning analysis of Llama 3.1 8B using LoRA (Hu et al., 2022) trained explicitly on the test set. We also conducted a separate experiment to test judge model variation, using seven different judges: Mistral Large, GPT-4o, GPT-4o mini, Mixtral 87B, Mistral 7B, along with the original Llama 3.1 8B and its newly fine-tuned variant. Each debate followed our structured protocol (Section 3) with a double round-robin format, aggregating model win counts as the primary metric. or benchmark standardization, we included our dataset of 5,500 debate transcripts—which collectively contain over 11,000 argumentative rounds—spanning the full double round-robin interactions among the eleven reference models, tested various scoring methods including Elo, Bradley–Terry, and TrueSkill, and ultimately selected TrueSkill due to its robustness on partial data.
@@ 5.1 Overall Performance Comparison
We further analyze debate performance by isolating results from models exclusively in the defender (Pro) role, reflecting their ability to justify and maintain correct solutions. Figure 4 compares total debate wins (combined Pro and Con) to Defend-Only Wins and Questioning-Only Wins, illustrating the consistency of model rankings across these roles.
@@ 5.2 Role-Specific Performance Analysis
The rankings based solely on Defend-Only Wins perfectly align with overall debate win rankings for the top nine models, underscoring the robustness of our evaluation framework. The minor discrepancy observed between GPT-3.5-turbo and Mistral 7B at the bottom is negligible, indicating that a model’s overall debate success consistently reflects its genuine understanding and problem-solving capabilities.
@@
(b) Defend-Only Win Rates
@@ 5.3 Head-to-Head Model Comparisons
Pairwise Heatmaps and Transitivity Analysis. We constructed pairwise win rate heatmaps across three scenarios—questioning, defending, and overall—to refine the rankings in Table 1. Stronger models consistently dominated weaker ones, typically securing 70–90% win rates. While a few minor loops were observed in the defend-only and question-only settings, notably, the combined heatmap resolved almost all of these inconsistencies. Out of 55 unique model pairings, the combined results produced a remarkably consistent hierarchy, with only a single, marginal transitivity violation. This near-perfect (98%+) transitivity strongly suggests our role-switching approach effectively mitigates positional bias and data contamination, where contaminated models may artificially excel in defense roles. This high degree of fidelity implies our evaluation genuinely captures meaningful differences in models’ reasoning capabilities. For instance, even when provided with the correct answer, the weaker model (Mistral 7B) achieved only a 10% win rate against DeepSeek V3, clearly underscoring our framework’s effectiveness in distinguishing true model competence. Furthermore, this high level of consistency was confirmed in our supplementary evaluation on the GPQA benchmark (Appendix C), where the rankings exhibited perfect transitivity, underscoring the robustness of the method across different datasets.
@@ 6 Discussion and Conclusion
Appendix A Additional Debate Evaluation: Question-Only Heatmap

### v2

@@ 1 Introduction
We introduce a systematic, reproducible evaluation pipeline transforming standard QA datasets into structured adversarial debates, effectively reducing subjective biases, clearly signaling model reasoning capabilities, and significantly decreasing benchmark curation costs. Empirical experiments with various commercial and open-source models on MMLU-Pro (Wang et al., 2024) validate our method’s robustness and resistance to data contamination. Additionally, weaker judges reliably distinguished stronger debaters, reinforcing findings by Khan et al. (2024), thereby ensuring scalability and future-proof applicability for advanced AI evaluations. We plan to publicly release a standardized benchmark derived from 5,500 structured debates—totaling over 11,000 rounds of argumentation—involving eleven reference models, providing the research community with a dependable evaluation resource. 111Code for this project is available at: https://github.com/l6cao/Debate-Driven-Evaluation.
@@ 4 Experimental Setup
To assess data contamination effects, we performed a fine-tuning analysis of Llama 3.1 8B using LoRA (Hu et al., 2022) trained explicitly on the test set. We also conducted a separate experiment to test judge model variation, using seven different judges: Mistral Large, GPT-4o, GPT-4o mini, Mixtral 87B, Mistral 7B, along with the original Llama 3.1 8B and its newly fine-tuned variant. Each debate followed our structured protocol (Section 3) with a double round-robin format, aggregating model win counts as the primary metric. For benchmark standardization, we included our dataset of 5,500 debate transcripts—which collectively contain over 11,000 argumentative rounds—spanning the full double round-robin interactions among the eleven reference models, tested various scoring methods including Elo, Bradley–Terry, and TrueSkill, and ultimately selected TrueSkill due to its robustness on partial data.
@@ 5.1 Overall Performance Comparison
We further analyze debate performance by isolating results from models exclusively in the defender (Pro) role, reflecting their ability to justify and maintain correct solutions. Figure 4 compares total debate wins (combined Pro and Con) to Defending-Only Wins and Questioning-Only Wins, illustrating the consistency of model rankings across these roles.
@@ 5.2 Role-Specific Performance Analysis
The rankings based solely on Defending-Only Wins perfectly align with overall debate win rankings for the top nine models, underscoring the robustness of our evaluation framework. The minor discrepancy observed between GPT-3.5-turbo and Mistral 7B at the bottom is negligible, indicating that a model’s overall debate success consistently reflects its genuine understanding and problem-solving capabilities.
@@
(b) Defending-Only Win Rates
@@ 5.3 Head-to-Head Model Comparisons
Pairwise Heatmaps and Transitivity Analysis. We constructed pairwise win rate heatmaps across three scenarios—questioning, defending, and overall—to refine the rankings in Table 1. Stronger models consistently dominated weaker ones, typically securing 70–90% win rates. While a few minor loops were observed in the defending-only and questioning-only settings, notably, the combined heatmap resolved almost all of these inconsistencies. Out of 55 unique model pairings, the combined results produced a remarkably consistent hierarchy, with only a single, marginal transitivity violation. This near-perfect (98%+) transitivity strongly suggests our role-switching approach effectively mitigates positional bias and data contamination, where contaminated models may artificially excel in defense roles. This high degree of fidelity implies our evaluation genuinely captures meaningful differences in models’ reasoning capabilities. For instance, even when provided with the correct answer, the weaker model (Mistral 7B) achieved only a 10% win rate against DeepSeek V3, clearly underscoring our framework’s effectiveness in distinguishing true model competence. Furthermore, this high level of consistency was confirmed in our supplementary evaluation on the GPQA benchmark (Appendix C), where the rankings exhibited perfect transitivity, underscoring the robustness of the method across different datasets.
@@ 6 Discussion and Conclusion
Appendix A Additional Debate Evaluation: Questioning-Only Heatmap